# Quanhao Li1∗  Zhen Xing1∗   Rui Wang1   Hui Zhang1  Qi Dai2  Zuxuan Wu1

### v1

@@ Abstract
Recent advances in video generation have led to remarkable improvements in visual quality and temporal coherence. Upon this, trajectory-controllable video generation has emerged to enable precise object motion control through explicitly defined spatial paths. However, existing methods struggle with complex object movements and multi-object motion control, resulting in imprecise trajectory adherence, poor object consistency, and compromised visual quality. Furthermore, these methods only support trajectory control in a single format, limiting their applicability in diverse scenarios. Additionally, there is no publicly available dataset or benchmark specifically tailored for trajectory-controllable video generation, hindering robust training and systematic evaluation. To address these challenges, we introduce MagicMotion, a novel image-to-video generation framework that enables trajectory control through three levels of conditions from dense to sparse: masks, bounding boxes, and sparse boxes. Given an input image and trajectories, MagicMotion seamlessly animates objects along defined trajectories while maintaining object consistency and visual quality. Furthermore, we present MagicData, a large-scale trajectory-controlled video dataset, along with an automated pipeline for annotation and filtering. We also introduce MagicBench, a comprehensive benchmark that assesses both video quality and trajectory control accuracy across different numbers of objects. Extensive experiments demonstrate that MagicMotion outperforms previous methods across various metrics. Our project page are publicly available at https://quanhaol.github.io/magicmotion-site/.
@@
equal contributions.
@@ 1 Introduction
With the rapid development of diffusion models, video generation has witnessed significant progress in recent years. Earlier video generation approaches, such as AnimateDiff [16] and SVD [3], primarily rely on the UNet [48] structure, which results in videos with limited length and quality. Sora [4] has demonstrated the powerful capabilities of the DiT [38] architecture in text-to-video (T2V) generation. Following this breakthrough, subsequent models leveraging the DiT architecture [80, 73, 11] have achieved higher quality outputs and extended video duration.
@@
We also construct MagicData, a high-quality public dataset comprising 51K video samples, each annotated with a <video, text, trajectory> triplet. We design a data pipeline that uses a large language model [53] to extract the main moving objects in the video, and Segment Anything Model (SAM2) [46, 45] to annotate the segmentation masks and bounding boxes of these objects. Furthermore, we introduce MagicBench, a large-scale comprehensive trajectory-controllable video generation benchmark. It categorizes all videos into 6 classes based on the number of foreground objects and evaluates models separately for each category in terms of both video generation quality and trajectory control accuracy.
@@ 2 Related Works
Video Diffusion Models   Diffusion models [19, 51, 52, 32] have made great progress in image generation [8, 37, 41, 47, 49, 42, 76], which has led to the rapid development of video generation [20, 63, 26, 16, 3, 22, 73, 68, 67, 54]. VDM [20] is the first to apply diffusion models to video generation. Early works like AnimateDiff [16] and SimDA [66] attempt to insert temporal layers into pretrained T2I model for video generation. Subsequently, VideoCrafter [6] and SVD [3] use large-scale and high-quality data for training, acheiving better performance. However, these methods have trouble on generating long videos with high quality, mainly due to the inherent limitations of the UNet architecture. The emergence of Sora [4] is a significant success, demonstrating the potential of DiT [38] models to generate high-quality videos with tens of seconds. Recent video generation methods [80, 73, 11] are mainly based on DiT architecture, and have achieved great success in the open-source community. However, these methods rely solely on text or image guidance for video generation, lacking precise control over object or camera trajectory, which is crucial for high-quality video generation.
Trajectory Controllable Video Generation   Trajectory Controllable Video Generation has recently garnered significant attention for its ability to precisely control object and camera trajectories during video synthesis. Previous methods [75, 78, 50, 31] integrate optical flow maps into video generation through a trajectory encoder. Recent works [59, 17] suggest using point maps as a form of guidance. MotionCtrl [59] processes point maps with a Gaussian filter and employs trainable encoders to encode object trajectories. Trackgo [17] represents objects using a few key points and injects this information via an encoder and a custom-designed adapter structure. Other works [56, 40, 36, 23] employ bounding box to control object trajectories. Boximator [56] employs a trainable self-attention layer to fuse box and visual tokens inspired by GLIGEN [30]. Some training-free methods [40, 36, 23] purpose to modify attention layers or initial noised video latents to inject box signals. Additionally, certain methods [57, 15, 13, 55] explore the potential of 3D trajectories to achieve more sophisticated motion control. LeViTor [55] employs keypoint trajectory maps enriched with depth information, while others [57, 15, 13] construct custom 3D trajectories to represent object movements. However, sparse trajectories lead to imprecise control on objects shape and size, while dense trajectories are difficult for users to provide. In contrast, MagicMotion can control both dense and sparse trajectories, providing users with more flexible control over video generation.
@@ 3.2 Model Architecture
We utilize CogVideoX-5B-I2V  [73] as our base image to video model. CogVideoX is built upon a DiT (Diffusion Transformer) architecture, incorporating 3D-Full Attention to generate high-quality videos. As shown in Fig. 2, the model takes an input image $I\in R^{H\times W\times 3}$ and a corresponding video $V\in R^{T\times H\times W\times 3}$ and encodes them into latent representations $Z_{image},Z_{video}\in R^{\frac{T}{4}\times\frac{H}{8}\times\frac{W}{8}\times 16}$ using a pretrained 3D VAE  [27]. Later, $Z_{image}$ is zero-padded to $T$ frames and concatenated with a noised version of $Z_{video}$ and then fed into the Diffusion Transformer, where a series of Transformer blocks iteratively denoise it over a predefined number of steps. Finally, the denoised latent is decoded by a 3D VAE decoder to get the output video $V_{out}\in R^{T\times H\times W\times 3}$.
@@ 3.3 Dense-to-Sparse Training Procedure
Dense trajectory conditions, such as segmentation masks, offer more precise control than sparse conditions like bounding boxes but are less user-friendly. To address this, MagicMotion employs a progressive training procedure, where each stage initializes its model with the weights from the previous stage. This enables three types of trajectory control ranging from dense to sparse. We found that this progressive training strategy helps the model achieve better performance compared to training from scratch with sparse conditions.
Specifically, we adopt the following trajectory conditions across stages: stage1 uses segmentation masks, stage2 uses bounding boxes, and stage3 uses sparse bounding boxes, where fewer than 10 frames have box annotations. Additionally, we always set the first frame of the trajectory condition as a segmentation mask to specify the foreground objects that should be moving.
@@ 3.4 Latent Segmentation Loss
To incorporate dense trajectory information while keeping computational costs within a reasonable range, we propose utilizing a lightweight segmentation head to predict segmentation masks directly in the latent space, eliminating the need for decoding operations.
Specifically, our segmentation head takes a list of diffusion features $Z_{feature}\in R^{\frac{T}{4}*\frac{H}{16}*\frac{W}{16}*3072}$ from each DiT block, and outputs a latent segmentation mask $Z_{segment}\in R^{\frac{T}{4}*\frac{H}{8}*\frac{W}{8}*16}$. We use a light-weight architecture inspired by Panoptic FPN [28]. Each diffusion feature first passes through a convolution layer to extract visual features $P\in R^{\frac{T}{4}*\frac{H}{16}*\frac{W}{16}*64}$. The resulting features are then concatenated and processed by another convolution layer followed by an upsampling layer to generate the final latent segmentation mask.
We compute the latent segment loss as the Euclidean distance between $Z_{segment}$ and the ground truth mask trajectory latents $Z_{mask}$, which can be written as,
@@ 3.5 Data Pipeline
Curation Pipeline.   We begin our dataset curation process with Pexels [24], a large-scale video-text dataset containing 396K video clips with text annotations. It encompasses videos featuring diverse subjects, various scenes, and a wide range of movements. We utilize Llama3.1 [53] to extract the foreground moving objects from the textual annotations of each video. As shown in Fig. 3, we input the video’s caption into the language model and prompt it to identify the main foreground objects mentioned in the sentence. If the model determines that the sentence does not contain any foreground objects, it simply returns “empty” and such videos are filtered out. Next, we utilize Grounded-SAM2 [44, 46], a grounded segmentation model that takes a video along with its main objects as input and generates segmentation masks for each primary object. Each object is consistently annotated with a unique color. Finally, bounding boxes are extracted from each segmentation mask using the coordinates of the top-left and bottom-right corners to draw the corresponding boxes. The color of the bounding box for each object remains consistent with its segmentation mask.
@@
Motion-I2V [50] 156.59 354.10 56.19 60.66 184.35 1558.32 32.00 42.61
ImageConductor [31] 234.35 331.76 51.76 52.90 224.32 1155.08 34.39 43.41
DragAnything [64] 137.57 253.40 66.30 70.85 180.77 1166.22 40.13 53.60
LeViTor [55] 189.63 194.53 39.96 46.36 225.42 922.68 25.24 31.42
DragNUWA [75] 140.99 185.52 66.88 69.21 182.70 1079.89 41.22 52.61
SG-I2V [36] 117.04 168.82 68.78 74.39 177.39 1170.60 37.36 50.96
Tora [78] 144.49 245.23 58.95 64.03 138.55 766.76 37.98 50.90
Ours (Stage1) 87.13 112.69 91.57 87.75 117.12 579.94 81.33 84.97
Ours (Stage2) 93.27 107.21 76.61 81.45 124.92 760.95 53.94 72.84
@@
Specifically, based on extensive manual evaluation, we empirically set the optical flow score threshold to 2.0, limit the number of foreground object annotations from 1 to 3, and constrain the annotated area ratio to a range of 0.008 to 0.83. The whole data curation and filtering pipeline yields us with MagicData, a high quality dataset for trajectory controllable video generation containing 51K videos with both dense and sparse trajectory annotations.
@@ 4.1 Experiment Settings
We employ CogVideoX 5B [73] as our base image-to-video model, which is trained to generate a 49-frame video at a resolution of $480\times 720$. Each stage of MagicMotion was trained on MagicData for one epoch. The training process consists of three stages. stage1 trains Trajectory ControlNet from scratch. In stage2, Trajectory ControlNet is further refined using the weights from stage1, while Segment Head is trained from scratch. Finally, in stage3, both Trajectory ControlNet and Segment Head continue training initialized with the weights from stage2. All training experiments were conducted on 4 NVIDIA A100-80G GPUs. We employed AdamW  [33] as the optimizer, training with a learning rate of $51e-$ 5 and a batch size of 1 on each GPU. During inference, we set steps to 50, the guidance scale to 6, and the weight of Trajectory ControlNet to 1.0 by default.
@@
During training, we use MagicData as our training set. MagicData is annotated with dense to sparse trajectory information using the data pipeline described in Section 3.5. It comprises a total of 51,000 <video, text, trajectory> triplets. During the training process, each video was resized to $480\times 720$ and sampled out 49 frames per clip. For evaluation, we evaluate all methods on both MagicBench and DAVIS [39], using the comparison metrics illustrated in Sec 3.6.
@@ 4.2 Comparison with Other Approaches
As shown in Table 1, our method outperforms all previous approaches across all metrics both on MagicBench and DAVIS, demonstrating its ability to produce higher-quality videos and more precise trajectory control. Additionally, we evaluate each method’s performance on MagicBench based on the number of controlled objects. As shown in Fig. 4, our method achieves the best results across all object number categories, further demonstrating the superiority of our approach.
@@ 4.3 Ablation Studies
Ablations on Latent Segment Loss.   Latent Segment Loss makes the model predict dense segmentation masks while training with sparse trajectories, enhancing its ability to perceive fine-grained object shapes under sparse conditions. To evaluate the effectiveness of this technique, we train the model from stage1 for one epoch using bounding boxes as trajectory conditions and compare its performance with MagicMotion stage2. Table 3 shows that the absence of Latent Segment Loss reduces the model’s ability on object shapes, leading to less precise trajectory control. Qualitative comparisons in Fig. 8 further highlight this effect. Without Latent Segment Loss, the woman’s arm in the generated video appears incomplete.
@@ 5 Conclusion
Appendix A Additional Experiments
@@
Appendix B Latent Segment Masks
@@
Appendix C Additional Comparison results
@@
MagicMotion 480*720 49 CogVideoX [73]
@@
Motion-I2V [50] 148.4379 660.8655 0.6057 0.7142 178.7882 867.6057 0.5078 0.5938
ImageConductor [31] 236.6626 674.4987 0.5706 0.6974 252.6367 879.8890 0.4786 0.5579
DragAnything [64] 147.2262 884.6453 0.6706 0.8088 169.6339 940.4941 0.6148 0.7232
LeViTor [55] 176.2211 492.4725 0.5536 0.7057 212.8940 542.8533 0.4352 0.5364
DragNUWA [75] 141.6972 610.0368 0.6699 0.7769 163.5423 624.3985 0.6033 0.6809
SG-I2V [36] 129.3680 547.8107 0.7144 0.8668 142.6613 619.2791 0.6315 0.7378
Tora [78] 124.8919 805.0145 0.6468 0.7776 156.2070 795.8535 0.5509 0.6584
Ours (Stage1) 96.3028 473.2179 0.9359 0.9607 111.3980 428.3430 0.9080 0.9097
Ours (Stage2) 106.8807 564.1036 0.7363 0.9017 122.6654 550.5857 0.6931 0.8256
@@
Motion-I2V [50] 168.2760 842.6530 0.5366 0.6076 149.3808 744.5470 0.6018 0.6484
ImageConductor [31] 243.8881 927.3884 0.5169 0.5578 221.8339 832.4498 0.5679 0.5417
DragAnything [64] 144.5718 925.2795 0.6332 0.6625 120.3439 901.7427 0.6946 0.7148
LeViTor [55] 195.1743 607.7522 0.3809 0.4671 183.5972 688.8164 0.3555 0.4044
DragNUWA [75] 145.0172 642.4184 0.6420 0.7012 122.8971 512.5130 0.7085 0.7250
SG-I2V [36] 126.4608 520.2733 0.6531 0.7068 97.3747 460.1303 0.7145 0.7423
Tora [78] 126.4421 742.4080 0.5926 0.6356 115.4566 779.0798 0.6226 0.6312
Ours (Stage1) 89.1128 421.0036 0.9107 0.8797 73.5877 396.4754 0.9231 0.8896
Ours (Stage2) 97.4697 440.2373 0.7562 0.8097 77.4146 442.0640 0.7998 0.8253
@@
Motion-I2V [50] 148.1927 582.0029 0.6295 0.6267 146.4851 923.2557 0.4899 0.4511
ImageConductor [31] 235.7079 857.3489 0.5180 0.4737 215.3864 963.6862 0.4536 0.3442
DragAnything [64] 120.7039 710.5812 0.7050 0.7011 122.6998 719.2442 0.6534 0.6045
LeViTor [55] 180.9895 578.5567 0.3913 0.4281 188.8038 763.3157 0.2812 0.2768
DragNUWA [75] 130.7921 435.9205 0.7253 0.6988 142.0180 549.7680 0.6638 0.5709
SG-I2V [36] 96.1895 453.1147 0.7431 0.7367 103.7038 596.4075 0.6616 0.6211
Tora [78] 117.2917 709.1618 0.6228 0.6111 148.9100 907.9254 0.4976 0.4866
Ours (Stage1) 76.4964 374.6467 0.9155 0.8600 75.8724 449.3122 0.9012 0.7653
Ours (Stage2) 79.5924 350.5010 0.8106 0.8123 75.6016 396.0661 0.8004 0.7124
@@
Appendix D Additional Ablation Results.
@@
Appendix E More Details on MagicData

### v3

# Quanhao Li1,2∗  Zhen Xing1,2∗   Rui Wang1,2   Hui Zhang1,2  Qi Dai3  Zuxuan Wu1,2†

@@ Abstract
Recent advances in video generation have led to remarkable improvements in visual quality and temporal coherence. Upon this, trajectory-controllable video generation has emerged to enable precise object motion control through explicitly defined spatial paths. However, existing methods struggle with complex object movements and multi-object motion control, resulting in imprecise trajectory adherence, poor object consistency, and compromised visual quality. Furthermore, these methods only support trajectory control in a single format, limiting their applicability in diverse scenarios. Additionally, there is no publicly available dataset or benchmark specifically tailored for trajectory-controllable video generation, hindering robust training and systematic evaluation. To address these challenges, we introduce MagicMotion, a novel image-to-video generation framework that enables trajectory control through three levels of conditions from dense to sparse: masks, bounding boxes, and sparse boxes. Given an input image and trajectories, MagicMotion seamlessly animates objects along defined trajectories while maintaining object consistency and visual quality. Furthermore, we present MagicData, a large-scale trajectory-controlled video dataset, along with an automated pipeline for annotation and filtering. We also introduce MagicBench, a comprehensive benchmark that assesses both video quality and trajectory control accuracy across different numbers of objects. Extensive experiments demonstrate that MagicMotion outperforms previous methods across various metrics.
@@
* equal contributions. $\dagger$} Corresponding author.
@@ 1 Introduction
With the rapid development of diffusion models, video generation has witnessed significant progress in recent years. In contrast to images, videos [74, 54, 64, 65] capture dynamic temporal changes and offer a richer representation of semantic content. Earlier video generation approaches, such as AnimateDiff [16] and SVD [3], primarily rely on the UNet [47] structure, which results in videos with limited length and quality. Sora [4] has demonstrated the powerful capabilities of the DiT [38] architecture in text-to-video (T2V) generation. Following this breakthrough, subsequent models leveraging the DiT architecture [94, 83, 28] have achieved higher quality outputs and extended video duration.
@@
We also construct MagicData, a high-quality public dataset comprising 23K video samples, each annotated with a <video, text, trajectory> triplet. We design a data pipeline that uses a large language model [53] to extract the main moving objects in the video, and Segment Anything Model (SAM2) [45, 44] to annotate the segmentation masks and bounding boxes of these objects. Furthermore, we introduce MagicBench, a large-scale comprehensive trajectory-controllable video generation benchmark. It categorizes all videos into 6 classes based on the number of foreground objects and evaluates models separately for each category in terms of both video generation quality and trajectory control accuracy.
@@ 2 Related Works
Video Diffusion Models
Diffusion models [18, 50, 51, 32] have made great progress in image generation [8, 37, 41, 46, 48, 42, 87, 89, 86, 88, 62, 91], which has led to the rapid development of video generation [19, 71, 25, 16, 3, 94, 21, 83, 28, 78, 77, 75, 73, 11, 55, 56, 57, 58, 52]. VDM [19] is the first to apply diffusion models to video generation. Early works like AnimateDiff [16] and SimDA [76] attempt to insert temporal layers into pretrained T2I model for video generation. Subsequently, VideoCrafter [6] and SVD [3] use large-scale and high-quality data for training, acheiving better performance. However, these methods have trouble on generating long videos with high quality, mainly due to the inherent limitations of the UNet architecture. The emergence of Sora [4] is a significant success, demonstrating the potential of DiT [38] models to generate high-quality videos with tens of seconds. Recent video generation methods [94, 83, 28] are mainly based on DiT architecture, and have achieved great success in the open-source community. However, these methods rely solely on text or image guidance for video generation, lacking precise control over object or camera trajectory, which is crucial for high-quality video generation.
Trajectory Controllable Video Generation
Trajectory Controllable Video Generation has recently garnered significant attention for its ability to precisely control object and camera trajectories during video synthesis. Previous methods [85, 92, 49, 31] integrate optical flow maps into video generation through a trajectory encoder. Recent works [67, 95] suggest using point maps as a form of guidance. MotionCtrl [67] processes point maps with a Gaussian filter and employs trainable encoders to encode object trajectories. Trackgo [95] represents objects using a few key points and injects this information via an encoder and a custom-designed adapter structure. Other works [61, 40, 36, 22] employ bounding box to control object trajectories. Boximator [61] employs a trainable self-attention layer to fuse box and visual tokens inspired by GLIGEN [30]. Some training-free methods [40, 36, 22] purpose to modify attention layers or initial noised video latents to inject box signals. Additionally, certain methods [63, 15, 13, 60] explore the potential of 3D trajectories to achieve more sophisticated motion control. LeViTor [60] employs keypoint trajectory maps enriched with depth information, while others [63, 15, 13] construct custom 3D trajectories to represent object movements. However, sparse trajectories lead to imprecise control on objects shape and size, while dense trajectories are difficult for users to provide. In contrast, MagicMotion can control both dense and sparse trajectories, providing users with more flexible control over video generation.
@@ 3.2 Model Architecture
We utilize CogVideoX-5B-I2V  [83] and Wan2.1 1.3B [59] as our base image to video model. The models are built upon a DiT (Diffusion Transformer) architecture, incorporating 3D-Full Attention to generate high-quality videos. As shown in Fig. 2, the model takes an input image $I\in R^{H\times W\times 3}$ and a corresponding video $V\in R^{T\times H\times W\times 3}$ and encodes them into latent representations $Z_{image},Z_{video}\in R^{\frac{T}{4}\times\frac{H}{8}\times\frac{W}{8}\times 16}$ using a pretrained 3D VAE  [26]. Later, $Z_{image}$ is zero-padded to $T$ frames and concatenated with a noised version of $Z_{video}$ and then fed into the Diffusion Transformer, where a series of Transformer blocks iteratively denoise it over a predefined number of steps. Finally, the denoised latent is decoded by a 3D VAE decoder to get the output video $V_{out}\in R^{T\times H\times W\times 3}$.
@@ 3.3 Dense-to-Sparse Training Procedure
Dense trajectory conditions, such as segmentation masks, offer more precise control than sparse conditions like bounding boxes but are less user-friendly. To address this, MagicMotion employs a progressive training procedure, where each stage initializes its model with the weights from the previous stage. This enables three types of trajectory control ranging from dense to sparse. We found that this progressive training strategy helps the model achieve better performance compared to training from scratch with sparse conditions. Specifically, we adopt the following trajectory conditions across stages: stage1 uses segmentation masks, stage2 uses bounding boxes, and stage3 uses sparse bounding boxes, where fewer than 10 frames have box annotations. Additionally, we always set the first frame of the trajectory condition as a segmentation mask to specify the foreground objects that should be moving.
@@ 3.4 Latent Segmentation Loss
To incorporate dense trajectory information while keeping computational costs within a reasonable range, we propose utilizing a lightweight segmentation head to predict segmentation masks directly in the latent space, eliminating the need for decoding operations. Specifically, our segmentation head takes a list of diffusion features $Z_{feature}\in R^{\frac{T}{4}*\frac{H}{16}*\frac{W}{16}*3072}$ from each DiT block, and outputs a latent segmentation mask $Z_{segment}\in R^{\frac{T}{4}*\frac{H}{8}*\frac{W}{8}*16}$. We use a light-weight architecture inspired by Panoptic FPN [27]. Each diffusion feature first passes through a convolution layer to extract visual features $P\in R^{\frac{T}{4}*\frac{H}{16}*\frac{W}{16}*64}$. The resulting features are then concatenated and processed by another convolution layer followed by an upsampling layer to generate the final latent segmentation mask. We compute the latent segment loss as the Euclidean distance between $Z_{segment}$ and the ground truth mask trajectory latents $Z_{mask}$, which can be written as,
@@ 3.5 Data Pipeline
Curation Pipeline.   We begin our dataset curation process with Pexels [23], a large-scale video-text dataset containing 396K video clips with text annotations. It encompasses videos featuring diverse subjects, various scenes, and a wide range of movements. We utilize Llama3.1 [53] to extract the foreground moving objects from the textual annotations of each video. As shown in Fig. 3, we input the video’s caption into the language model and prompt it to identify the main foreground objects mentioned in the sentence. If the model determines that the sentence does not contain any foreground objects, it simply returns “empty” and such videos are filtered out. Next, we utilize Grounded-SAM2 [44, 45], a grounded segmentation model that takes a video along with its main objects as input and generates segmentation masks for each primary object. Each object is consistently annotated with a unique color. Finally, bounding boxes are extracted from each segmentation mask using the coordinates of the top-left and bottom-right corners to draw the corresponding boxes. The color of the bounding box for each object remains consistent with its masks.
Specifically, based on extensive manual evaluation, we empirically set the optical flow score threshold to 2.0, limit the number of foreground object annotations from 1 to 3, and constrain the annotated area ratio to a range of 0.008 to 0.83. The whole data curation and filtering pipeline yields us with MagicData, a high quality dataset for trajectory controllable video generation containing 23K videos with both dense and sparse trajectory annotations.
@@
Motion-I2V [49] 39.31 354.10 56.19 60.66 94.14 1558.32 32.00 42.61
ImageConductor [31] 49.95 331.76 51.76 52.90 91.71 1155.08 34.39 43.41
DragAnything [72] 31.36 253.40 66.30 70.85 70.70 1166.22 40.13 53.60
LeViTor [60] 38.32 194.53 39.96 46.36 97.98 922.68 25.24 31.42
DragNUWA [85] 39.73 185.52 66.88 69.21 84.61 1079.89 41.22 52.61
SG-I2V [36] 32.60 168.82 68.78 74.39 90.93 1170.60 37.36 50.96
Tora [92] 26.27 245.23 58.95 64.03 51.75 766.76 37.98 50.90
Ours (Stage1-Wan1.3B) 17.74 129.79 83.79 80.96 46.65 548.56 42.83 53.67
Ours (Stage2-Wan1.3B) 19.33 143.07 68.89 71.74 48.33 662.82 44.75 61.88
Ours (Stage1-CogVideoX) 15.06 112.69 91.57 87.75 45.06 579.94 81.33 84.97
Ours (Stage2-CogVideoX) 15.17 107.21 76.61 81.45 50.36 760.95 53.94 72.84
@@ 4.1 Experiment Settings
We employ CogVideoX 5B [83] and Wan2.1 1.3B [59] as our base image-to-video model, which is trained to generate video based on an input image. Each stage of MagicMotion was trained on MagicData for one epoch. The training process consists of three stages. stage1 trains Trajectory ControlNet from scratch. In stage2, Trajectory ControlNet is further refined using the weights from stage1, while Segment Head is trained from scratch. Finally, in stage3, both Trajectory ControlNet and Segment Head continue training initialized with the weights from stage2. All training experiments were conducted on 4 NVIDIA A100-80G GPUs. We employed AdamW  [33] as the optimizer, training with a learning rate of 1e51e-5 and a batch size of 1 on each GPU. During inference, we set steps to 50, the guidance scale to 6, and the weight of Trajectory ControlNet to 1.0 by default.
@@
During training, we use MagicData as our training set. MagicData is annotated with dense to sparse trajectory information using the data pipeline described in Section 3.5. It comprises a total of 23K video samples, annotated with text and trajectory. For evaluation, we evaluate all methods on both MagicBench and DAVIS [39], using the comparison metrics illustrated in Sec 3.6.
@@ 4.2 Comparison with Other Approaches
As shown in Table 1, our method outperforms all previous approaches across all metrics both on MagicBench and DAVIS, demonstrating its ability to produce higher-quality videos and more precise trajectory control. Additionally, we evaluate each method’s performance on MagicBench based on the number of controlled objects. As shown in Fig. 4, our method achieves the best results across all object number categories, demonstrating the superiority of our approach.
@@ 4.3 Ablation Studies
Ablations on Latent Segment Loss.   Latent Segment Loss makes the model predict dense segmentation masks while training with sparse trajectories, enhancing its ability to perceive fine-grained object shapes under sparse conditions. To evaluate the effectiveness of this technique, we train the model from stage1 for one epoch using bounding boxes as trajectory conditions and compare its performance with MagicMotion stage2. Table 3 shows that the absence of Latent Segment Loss reduces the model’s ability on object shapes, leading to less precise trajectory control.
@@ 5 Conclusion
Acknowledge This work was supported in part by the National Natural Science Foundation of China (Grant 62032006 and Grant 62472098).
Appendix A Ablations on Latent Segment Loss
We present qualitative results from the ablation study on Latent Segment Loss. As shown in Fig. 8, removing the Latent Segment Loss significantly reduces the model’s ability to capture object shapes, resulting in less accurate trajectory control. For instance, without this loss, the woman’s arm in the generated video appears incomplete.
Figure 9: Camera motion controlled results. By setting specific trajectory conditions, MagicMotion can control camera movements.
Appendix B Additional Experiments
@@
Appendix C Latent Segment Masks
@@
Appendix D Additional Comparison results
@@
MagicMotion-CogVideoX 480*720 49 CogVideoX [83]
MagicMotion-Wan1.3B 480*832 81 Wan [59]
@@
Motion-I2V [49] 78.3245 660.8655 0.6057 0.7142 90.4269 867.6057 0.5078 0.5938
ImageConductor [31] 106.1569 674.4987 0.5706 0.6974 109.2959 879.8890 0.4786 0.5579
DragAnything [72] 72.2494 884.6453 0.6706 0.8088 85.6511 940.4941 0.6148 0.7232
LeViTor [60] 80.9714 492.4725 0.5536 0.7057 94.0106 542.8533 0.4352 0.5364
DragNUWA [85] 76.3963 610.0368 0.6699 0.7769 81.2921 624.3985 0.6033 0.6809
SG-I2V [36] 60.8617 547.8107 0.7144 0.8668 71.8699 619.2791 0.6315 0.7378
Tora [92] 60.2737 805.0145 0.6468 0.7776 73.6499 795.8535 0.5509 0.6584
Ours (Stage1-Wan1.3B) 48.4211 461.1011 0.8518 0.9186 49.3071 438.2138 0.8143 0.8161
Ours (Stage2-Wan1.3B) 51.2964 555.1390 0.6693 0.8163 55.5581 569.0549 0.6318 0.7454
Ours (Stage1-CogVideoX) 42.3523 473.2179 0.9359 0.9607 48.6525 428.3430 0.9080 0.9097
Ours (Stage2-CogVideoX) 46.1080 564.1036 0.7363 0.9017 52.7296 550.5857 0.6931 0.8256
@@
Motion-I2V [49] 89.3195 842.6530 0.5366 0.6076 83.2024 744.5470 0.6018 0.6484
ImageConductor [31] 114.7662 927.3884 0.5169 0.5578 110.8037 832.4498 0.5679 0.5417
DragAnything [72] 85.1437 925.2795 0.6332 0.6625 71.8865 901.7427 0.6946 0.7148
LeViTor [60] 93.3111 607.7522 0.3809 0.4671 90.4561 688.8164 0.3555 0.4044
DragNUWA [85] 79.3941 642.4184 0.6420 0.7012 71.0587 512.5130 0.7085 0.7250
SG-I2V [36] 70.0600 520.2733 0.6531 0.7068 56.9318 460.1303 0.7145 0.7423
Tora [92] 66.6571 742.4080 0.5926 0.6356 64.0669 779.0798 0.6226 0.6312
Ours (Stage1-Wan1.3B) 51.7821 458.4122 0.8236 0.8122 47.8972 463.4390 0.8598 0.8288
Ours (Stage2-Wan1.3B) 56.9454 568.4777 0.6632 0.7089 50.0364 548.9445 0.7112 0.7252
Ours (Stage1) 42.9636 421.0036 0.9107 0.8797 37.6524 396.4754 0.9231 0.8896
Ours (Stage2) 45.4721 440.2373 0.7562 0.8097 37.4172 442.0640 0.7998 0.8253
@@
Motion-I2V [49] 84.2231 582.0029 0.6295 0.6267 79.9077 923.2557 0.4899 0.4511
ImageConductor [31] 117.8054 857.3489 0.5180 0.4737 106.3168 963.6862 0.4536 0.3442
DragAnything [72] 79.7015 710.5812 0.7050 0.7011 87.1509 719.2442 0.6534 0.6045
LeViTor [60] 90.8332 578.5567 0.3913 0.4281 92.5265 763.3157 0.2812 0.2768
DragNUWA [85] 77.2094 435.9205 0.7253 0.6988 78.5956 549.7680 0.6638 0.5709
SG-I2V [36] 65.2560 453.1147 0.7431 0.7367 95.7569 596.4075 0.6616 0.6211
Tora [92] 67.3827 709.1618 0.6228 0.6111 77.2571 907.9254 0.4976 0.4866
Ours (Stage1-Wan1.3B) 40.8559 384.3586 0.8645 0.8112 39.6606 395.2139 0.8136 0.6726
Ours (Stage2-Wan1.3B) 46.7040 491.7079 0.7461 0.7210 42.0005 454.2931 0.6941 0.5861
Ours (Stage1) 39.4636 374.6467 0.9155 0.8600 39.2044 449.3122 0.9012 0.7653
Ours (Stage2) 40.0656 350.5010 0.8106 0.8123 37.1917 396.0661 0.8004 0.7124
@@
Appendix E Additional Ablation Results.
@@
Appendix F More Details on MagicData